OpenStudy (anonymous):

The number of girls in Shana's class is 2 more than the number of boys. Four times the number of boys is seven more than three times the number of girls. How many girls are in Shana's class? A) 9 B) 11 C) 13 D) 15

OpenStudy (anonymous):

g=2+b 4b=7+3g 2 equations we can substitute the value of g 4b=7+3g 4b= 7+ 3(2+b) 4b=7+6+3b 4b-3b=13 b=13 g=2+b g=2+13 g=15

OpenStudy (anonymous):

This is fantastic! This helps a lot, and this showed us the process for some other problems. Thank you~